OpenStudy (sleepyjess):

Find the exact value of \(tan (arcsin (\dfrac{2}{5})).\)

OpenStudy (sleepyjess):

@abhisar

OpenStudy (anonymous):

draw a triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of an angle whose sine is \(\frac{2}{5}\) all you need is the third side, which you find via pythagoras

OpenStudy (michele_laino):

to which quadrant belongs the unknow arc, whose sin is 2/5?

OpenStudy (anonymous):

it is \[x=\sqrt{5^2-2^2}=\sqrt{25-4}=\sqrt{21}\]

OpenStudy (anonymous):

you want the tangent, which is "opposite over adjacent' making your answer \[\tan(\arcsin(\frac{2}{5}))=\frac{2}{\sqrt{21}}\]

OpenStudy (sleepyjess):

Thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@Michele_Laino it makes no difference what quadrant you are in

OpenStudy (michele_laino):

why? If you are in the first quadrant tan is positive, whereas if you are in second quadrant tan is negative

OpenStudy (anonymous):

reason is that \[\arcsin(\theta)\] is a number between \(-\frac{\pi}{2}\) and \(\frac{\pi}{2}\)

OpenStudy (michele_laino):

@satellite73 Sorry, equation sin x=2/5 have two mainly solutions 

OpenStudy (anonymous):

ok i said that wrong the arcsine of a number cannot be in the second quadrant is what i meant to say

OpenStudy (anonymous):

oh no, you do not have two solutions, you have one solution only arcsine is a well defined function

OpenStudy (anonymous):

you are right in that \[\sin(x)=\frac{2}{5}\] has more than one solution in fact it has an infinite number of solutions

OpenStudy (anonymous):

but there is only one \[\arcsin(\frac{2}{5})\] it is the solution to \[\sin(x)=\frac{2}{5}\] only for \[-\frac{\pi}{2}\le x\le\frac{\pi}{2}\]

OpenStudy (michele_laino):

@satellite73 sorry, note that I said mainly solutions, If x is a solution, also 180-x is a solution

OpenStudy (anonymous):

like i said you have an infinite number of solutions to \[\sin(x)=\frac{2}{5}\] that is true but only one number is \[\arcsin(\frac{2}{2})\] since arcsine is a function (cannot have two answers)

OpenStudy (anonymous):

for example \[\arcsin(\frac{1}{2})=\frac{\pi}{6}\] not \(\frac{\pi}{6}\) or \(\frac{5\pi}{6}\)

OpenStudy (michele_laino):

@satellite73 Ok! you are right! sorry!

OpenStudy (anonymous):

no need to be sorry, it is common concern if the question was "find the tangent of a number whose sine is 2/5" then you would have been right, you would not know if it was positive or negative

OpenStudy (michele_laino):

@satellite73 thank you! :)